sam is 4 years older than cathy. if 4 times catherine's age is subtracted from 5 times sam's age, the difference is 32 years. find the age of each.
Answer by TutorDelphia(193)

sam is 4 years older than cathy.
s=c+4
4 times catherine's age is subtracted from 5 times sam's age, the difference is 32 years
5s-4c=32
and then I can solve using substitution
5(c+4)-4c=32
5c+20-4c=32
c+20=32
c=12
Sub in 12 for c
s=12+4
s=16